# CMOB: Large-Scale Cancer Multi-Omics Benchmark with Open Datasets, Tasks, and Baselines

### v1

@@ Abstract
Machine learning has shown great potential in the field of cancer multi-omics studies, offering incredible opportunities for advancing precision medicine. However, the challenges associated with dataset curation and task formulation pose significant hurdles, especially for researchers lacking a biomedical background. Here, we introduce the CMOB, the first large-scale cancer multi-omics benchmark integrates the TCGA platform, making data resources accessible and usable for machine learning researchers without significant preparation and expertise.To date, CMOB includes a collection of 20 cancer multi-omics datasets covering 32 cancers, accompanied by a systematic data processing pipeline. CMOB provides well-processed dataset versions to support 20 meaningful tasks in four studies, with a collection of benchmarks. We also integrate CMOB with two complementary resources and various biological tools to explore broader research avenues.All resources are open-accessible with user-friendly and compatible integration scripts that enable non-experts to easily incorporate this complementary information for various tasks. We conduct extensive experiments on selected datasets to offer recommendations on suitable machine learning baselines for specific applications. Through CMOB, we aim to facilitate algorithmic advances and hasten the development, validation, and clinical translation of machine-learning models for personalized cancer treatments. CMOB is available on GitHub (https://github.com/chenzRG/Cancer-Multi-Omics-Benchmark).
1 Introduction
Cancer is a major public health concern with increasing incidence and leading to mortality. The National Cancer Institute (NCI) reports that the high costs of cancer care have been projected to grow to \$246.6 billion by 2030 (COS, 2023). Without further investment in research and prevention, cancer will cost the global economy an accumulated Int\$25.2 trillion in healthcare costs over the next 30 years (Kreier, 2023). Formerly, cancer research remains at the histomorphologic level. Cancer was classified solely according to organs of the origin or simplistic appearance features (Schneider et al., 2005). It is impossible to precisely distinguish and treat many morphologically similar and organ-homologous cancers. Over the past 15 years, increasing molecular and tumor biology knowledge has proven that cancer is a multi-omics disease (Hasin et al., 2017; Olivier et al., 2019; Vasaikar et al., 2018). Cancer is influenced by events occurring at various omics levels, which are interconnected (Balmain et al., 2003; Sandoval and Esteller, 2012; Jones and Baylin, 2007). For example, changes in genome sequences (genomics) regulate gene expressions (transcriptomics), finally reflecting on the phenotype, such as cancer subtypes, responses to treatment, and prognosis (Schneider and Orchard, 2011; Mani et al., 2022). This shift toward a multi-omics perspective led to a new era of cancer research.
Machine learning has proven helpful for mining multi-omics data (Reel et al., 2021a; Picard et al., 2021). However, the development of effective approaches for studying cancer multi-omics (or omics) remains relatively limited. This issue is quite apparent compared with the more common single-cell omics studies, which are currently the main focus of machine learning scientists (Wang and Bodovitz, 2010). The main reason is the lack of available data sources. Single-cell omics data can be easily generated in controlled in-lab experiments, and tens of thousands of cell-level cultured samples are available for training machine learning models (Lee et al., 2020). Cancer omics data come from clinical environments with data collection restrictions such as patient privacy. Each research project typically involves only a limited number of patient samples (Menyhárt and Győrffy, 2021b). Actually, a large amount of cancer omics data still belongs to individual research groups, hindering scientists from formulating cancer multi-omics studies as solvable machine-learning tasks.
There is an increasing effort to make cancer omics data publicly available. The Cancer Genome Atlas (TCGA) (The Cancer Genome Atlas Research Network, 2013) is the most established and well-used public multi-omics data resource. TCGA has molecularly characterized patient samples from various cancer types, providing the research community over 2.5 petabytes of multi-omics data. Despite its potential, the usability of TCGA presents challenges, particularly for researchers without a biomedical background (Tomczak et al., 2015).
(1) Lacking omics-ready datasets. While TCGA offers many cancer samples, they are not specifically tailored for omics studies and lack cross-omics alignment. Omics data processing involves several complex steps, such as categorizing omics, sample transformation, uniforming, or quality control, all of which require domain-specific expertise and bioinformatics tools. Therefore, it is necessary to establish a benchmark of cancer multi-omics datasets that can be used directly, even by non-experts.
(2) Lacking task-ready datasets. The fundamental unit in TCGA is the patient, which lacks precise categorization for specific cancer multi-omics tasks. Satisfied samples are often scattered across repositories and presented in complex, machine-unreadable file formats (Kircher and Kelso, 2010). Extensive curation is required to link samples with metadata and convert them into formats suitable for model analysis. Making data resources ready to learn computationally on cancer multi-omics tasks is very important.
(3) Non-compatibility with cross-platform databases. Studying cancer multi-omics often requires integration with biological corpora to maintain biological accuracy. For example, transcriptomics data can be mapped to graph-structured gene corpora (e.g., STRING (Szklarczyk et al., 2023) or GEO (Barrett et al., 2012)) to study gene regulation networks or Protein-Protein Interactions (PPI). However, combining TCGA with interaction-focused databases requires bioinformatics expertise, manual retrieval, and scripting skills to ensure accurate data integration. Standardizing data formats and providing seamless data integration could enhance the development of interdisciplinary machine-learning approaches.
To address the challenges outlined above, we introduce CMOB, the first Cancer Multi-Omics Benchmark that integrate TCGA platform and make data resource more accessible and usable for machine learning researchers without significant preparation and expertise. To summarize, our contributions are as follows:
  1. 1.
    Omics-ready datasets. CMOB includes carefully processed 20 datasets covering 32 cancer types with aligned patient samples across four major omics types, i.e., mRNA expression, microRNA expression, DNA methylation, and copy number variations. We assemble multiple feature scale versions for each dataset to adapt to downstream tasks. We also provide complete data processing pipelines to facilitate replication and user customization.
  2. 2.
    Tasks-ready datasets and baselines. CMOB offers 20 learning tasks in four studies, including cancer subtype identification, pan-cancer classification, and TCGA omics data imputation, each with a corresponding dataset version. For all benchmark tasks, we collect various baseline models from both the biomedical community and recent machine learning advancements. Moreover, all models are reproduced, evaluated, and extended under our benchmark datasets.
  3. 3.
    Complementary resources. To date, CMOB includes three complementary resources: the cross-platform STRING corpus, clinical health records, and downstream analysis tools. We provide user-friendly and compatible integration scripts that enable non-experts to easily incorporate this complementary information for broader biological research and clinical support and help researchers conduct classic and reliable biological validations of their data analysis frameworks.
Remark: There is a growing body of literature studying cancer omics in machine learning (Withnell et al., 2021; Chen et al., 2022; Yang et al., 2023). The results typically involve self-collected datasets from TCGA Silva et al. (2016). Some researchers make their data collection and processing source code publicly available. However, using these resources can be cumbersome and requires substantial biological knowledge and computational tools like R packages Zhu et al. (2014). Moreover, much of this code is scattered across various cancer research articles in the biomedical community, complicating accessibility for machine learning researchers Cline et al. (2013); Liu et al. (2018). The lack of comprehensive, publicly available databases hinders the progress in cancer omics studies.
The remainder of the paper proceeds as follows. After reviewing related work in Section 2, Section 3 introduces the construction of CMOB. Before concluding, in Section 4, we conduct the baseline experiments on selected datasets and report the results.
Figure 1: CMOB provides (1) 20 cancer multi-omics datasets covering 32 cancers and a systematic data processing pipeline; (2) well-processed dataset versions to support 20 meaningful tasks with a collection of benchmarks; and (3) complementary resources to connect cross-platform databases and support fundamental cancer biology or clinical studies.
2 Related Work
CMOB is the first benchmark that includes a variety of available datasets and learning tasks for cancer multi-omics data analysis. In this section, we discuss how CMOB integrates with machine learning methods for cancer omics studies and compare it with benchmarks in related areas.
2.1 Relation to Machine-Learning Methods for Cancer Omics Study
Over the past ten years, studying cancer omics has become a cornerstone of modern cancer research Chakraborty et al. (2018); Berg et al. (2017). Key tasks in omics data analysis include identifying clinical phenotypes such as tumor types and subtypes, drug responses, and survival outcomes Yang et al. (2021a); Liu et al. (2010); Menyhárt and Győrffy (2021a); Li et al. (2021a); Wang et al. (2020); Chen et al. (2023). Researchers are dedicated to developing computational methods to accelerate all aspects of cancer omics development Kaur et al. (2021); Chen et al. (2022); Arjmand et al. (2022). Earlier studies use some statistical-based methods from the biomedical community Reel et al. (2021b). For example, iClusterBayes Mo et al. (2018) integrates multi-omics data using a Bayesian latent variable regression model for downstream prediction tasks; SNF Wang et al. (2014) constructs sample-sample similarity networks at omics levels for cancer clustering. Additionally, there are notable deep learning-based approaches from the machine learning community, such as Subtype-GAN Yang et al. (2021b), which employs a generative adversarial network (GAN) structure to extract features from each omics source for cancer subtype identification. Other works propose using variational autoencoders (VAE) (Withnell et al., 2021) and Vector Quantized-VAE (Chen et al., 2022, 2023) to capture a meaningful latent space for distinguishing cancer samples. Recent attempts in contrastive learning have shown the effectiveness of integrating multi-omics data into a unified feature space (Yang et al., 2023). Cancer omics data are also pivotal in related areas like gene network analysis Chou et al. (2014); Heo et al. (2021). For example, Saha et al. Saha et al. (2017) utilized gene expression data to infer transcriptome-wide gene networks, while Aibar et al. Aibar et al. (2017) developed a workflow to map gene regulatory networks in omics data and identify stable cell states.
2.2 Relation to Other Benchmarks on Biomedical and Healthcare Data
Compound chemical structure data is one of the most discussed data forms in the biochemical community. Several benchmarks have been proposed to advance representation learning for compounds and proteins. For example, MoleculeNet offers datasets for molecular modeling, while TAPE Rao et al. (2019) provides tasks aimed at protein transfer learning. More relevant to our focus, although primarily concentrated on drug discovery and development, TDC Huang et al. (2021) encompasses a broader range of data modalities relevant to therapeutics, including compounds, biomolecular interactions, and genomic sequences. Moreover, there are several review studies Watson et al. (2022); Colomé-Tatché and Theis (2018); Miao et al. (2021); Ma et al. (2020) on single-cell omics data integration methods within bioinformatics and omics research. However, cancer multi-omics data is relatively absent in the data science and machine learning communities as an important data form in biomedical and healthcare. We address this gap by providing cancer multi-omics datasets benchmark, meaningful learning tasks, and extensive evaluations, positioning CMOB as a critical component in biomedical and healthcare data benchmarks.
3 Cancer Multi-Omics Benchmark (CMOB)
CMOB has three major components: $\rm(\hskip 1.79993pti\hskip 1.79993pt)$ a collection of cancer multi-omics datasets and a systematic data processing pipeline; $\rm(\hskip 1.79993ptii\hskip 1.79993pt)$ well-processed dataset versions to support meaningful tasks with a collection of benchmarks; and $\rm(\hskip 1.79993ptiii\hskip 1.79993pt)$ complementary resources to connect cross-platform databases and support fundamental cancer biology or clinical studies. An overview can be found in Figure 1.
3.1 Overall Datasets Composition
Table 1: CMOB provides a collection of 20 multi-omics datasets covering 32 cancer types, that is, one Pan-cancer dataset, nine unlabeled cancer subtype datasets, five labeled, golden-standard subtype datasets, and five TCGA data imputation datasets.
Dataset Feature Scale Omics Feature Size Sample Size #Baselines Learning Task
mRNA miRNA Methy CNV
ACC Orignal 18034 368 19045 19525 177 10 Clustering
KIRP Orignal 18465 769 18715 19551 273 10 Clustering
KIRC Orignal 18464 352 19045 19523 314 10 Clustering
LIHC Orignal 17946 846 18714 19551 364 10 Clustering
LUAD Orignal 18310 427 19052 19551 450 10 Clustering
LUSC Orignal 18206 423 19060 19551 363 10 Clustering
PRAD Orignal 17954 759 19049 19568 450 10 Clustering
THCA Orignal 17261 375 19052 19551 291 10 Clustering
THYM Orignal 18354 1018 18716 19551 119 10 Clustering
Pan-cancer Aligned 3217 383 3139 3105 8314 10 Classification
GS-COAD Orignal 17261 375 19052 19551 260 10 Classification
GS-BRCA Orignal 18206 368 19049 19568 671 10 Classification
GS-GBM Orignal 20684 335 19034 19545 243 10 Classification
GS-LGG Orignal 18345 345 19023 19534 246 10 Classification
GS-OV Orignal 17354 244 19034 19534 284 10 Classification
Imp-COAD Top 2000 200 2000 2000 260 7 Imputation
Imp-BRCA Top 2000 200 2000 2000 671 7 Imputation
Imp-GBM Top 2000 200 2000 2000 243 7 Imputation
Imp-LGG Top 2000 200 2000 2000 246 7 Imputation
Imp-OV Top 2000 200 2000 2000 284 7 Imputation
As shown in Table 1, CMOB provides a collection of 20 multi-omics datasets including:
-) One pan-cancer dataset involves patients with 32 cancer types.
-) Nine unlabeled cancer subtype datasets include Adrenocortical Carcinoma (ACC), Kidney Renal Papillary Cell Carcinoma (KIRP), Kidney Renal Clear Cell Carcinoma (KIRC), Liver Hepatocellular Carcinoma (LIHC), Lung Adenocarcinoma (LUAD), Lung Squamous Cell Carcinoma (LUSC), Prostate Adenocarcinoma (PRAD), Thyroid Carcinoma (THCA), and Thymoma (THY).
-) Five labeled, golden-standard subtype datasets corresponding to five cancers include Colon Adenocarcinoma (GS-COAD), Breast invasive carcinoma (GS-BRCA), Glioblastoma Multiforme (GS-GBM), Brain Lower Grade Glioma (GS-LGG), and Ovarian Serous Cystadenocarcinoma (GS-OV).
-) Five TCGA data imputation datasets include corrupted omics profiles from the above well-studied cancer types involving Imp-COAD, Imp-BRCA, Imp-GBM, Imp-LGG, and Imp-OV.
-) Two complementary data resources include a collected corpus from STRING and a collection of Electronic Health Records (EHR) data for cancer samples, accompanied by interactive scripts for integration.
Multiple-Scaled Feature. Cancer multi-omics analysis always suffers from an unbalanced sample and feature size. CMOB hence provides three versions of feature scales, i.e., Original, Top, and Aligned, to support feasible analysis. Original features are extracted directly from each dataset and correspond to the complete set of features without filtering. Users can customize their datasets. Top features are identified through ANOVA St et al. (1989) statistical testing according to p-values, selecting the most significant features among samples. This approach unifies the feature size and potentially reduces the noise features. Aligned features are determined by the intersection of features present across all sub-datasets, corresponding to the shared features among different sub-datasets.
3.2 Omics-Ready Datasets and Assembly Pipeline
Motivation. TCGA was initially established to support broad cancer research. It is not specifically tailored for multi-omics studies but collects data at each cancer genome characterization workflow step. Collecting multi-omics data from TCGA presents several challenges:
  • No omics-specific data were given. Raw data are in many different formats, and other resource enrollments, e.g., diagnostic imaging and clinical records, are also included.
  • No unified omics profiles (feature resources) were given. Omics profiles collected for one sample are scattered across repositories.
  • Sample retrieval, profile alignment, and data cleaning by bioinformatics tools are necessary to convert raw data into omics profiles suitable for analysis.
CMOB has removed these barriers with the assembled pipeline below:
(1) Omics Recruitment and Resource Collection: CMOB recruits four representative omics types that are most discussed in current cancer studies, that is, mRNA expression (mRNA), miRNA expression (miRNA), DNA methylation (Methy), and copy number variations (CNV). More detailed information on these omics types can be found in the Supplementary Material Section B. The raw omics data are obtained from TCGA through the Genomic Data Commons (GDC) Data Portal Grossman et al. (2016). While samples are provided by different biologically experimental platforms, we incorporate data from as many platforms as possible to include a broader range of patient samples. For instance, for Methy omics data, we chose platforms like "Illumina Human Methylation 27", "Illumina Human Methylation 450", and "Illumina Human Methylation EPIC". We verify the integrity of the files using checksums (as specified in the MANIFEST file) to ensure no files are corrupted or missing. The collected data files are then organized by cancer type for further processing.
(2) Omics-specific Data Formatting and Profiling: Each type of omics data requires a specific series of processing steps tailored to its unique biological characteristics. Below, we outline a general pipeline that summarizes these processes:
  • Omics Identification and Categorization. Identify the type of omics data from the metadata and ensure proper categorization (e.g., "Transcriptome Profiling" for transcriptomics).
  • Platform and Experimental Protocol Verification. Identify and confirm the experimental platform and workflow type from metadata (e.g., Illumina Hi-Seq for sequencing platforms).
  • Feature Alignment and Profiling. Convert raw data into omics features based on technical details in metadata. For example, aligning scaled RSEM estimates to FPKM for platforms like Illumina Hi-Seq. Key notions can be found in guidelines like "Average methylation (beta-values) of promoters defined as 500bp upstream & 50 downstream of Transcription Start Site (TSS)," which shows the gene methylation definitions.
  • Mensurment Transformation and Adjustments. Apply necessary transformations (e.g., logarithmic for transcriptomics or median-centering for methylation) to ensure data consistency.
  • Annotation and Saving. Annotate processed data to ensure that omics data is labeled with unified gene IDs. Store annotated data in appropriate categories based on the specific tasks.
More details on the data processing pipeline can be found in the Appendix Section B.
3.3 Task-Ready Datasets and Collected Baselines
As introduced in Section 3.1, CMOB currently provides 20 learning tasks in three studies, including pan-cancer classification, cancer subtype identification, and omics data imputation, each with a corresponding dataset version, baseline methods, and evaluation metrics.
(1) Pan-cancer Classification:
Motivation.This task aims to identify the specific cancer type for each patient, enhancing early diagnostic accuracy and potentially improving treatment outcomes.
Baseline Methods. Several computational multi-omics data integration methods have been proposed for cancer identification using classical statistical machine learning and deep-based methods. Currently, we have enrolled well-used, open-sourced statistical methods, including: Similarity Network Fusion (SNF) (Wang et al., 2014) integrates omics data by iteratively refining sample similarity networks and applying spectral clustering; Neighborhood-based Multi-Omics clustering (NEMO) (Rappoport and Shamir, 2019) converts sample similarity networks to relative similarity for group comparability; Cancer Integration via Multi-kernel Learning (CIMLR) (Wilson et al., 2019) combines various Gaussian kernels into a similarity matrix for clustering; iClusterBayes (Mo et al., 2018) projects input into a low-dimensional space using the Bayesian latent variable regression model for clustering; moCluster (Meng et al., 2016) uses multiple multivariate analyses to calculate latent variables for classification; We also enrolled more recent deep-dased methods, including: Subtype-GAN (Yang et al., 2021b) extracts features from each omics data by relatively independent GAN layers and integrates them; DCAP (Chai et al., 2021) integrates multi-omics data by the denoising autoencoder to obtain the representative features; MAUI (Ronen et al., 2019) uses stacked VAE to extract many latent factors to identify patient groups; XOmiVAE (Withnell et al., 2021) uses VAE for low-dimensional latent space extraction and classification; MCluster-VAEs (Rong et al., 2022) uses VAE with an attention mechanism to model multi-omics data.
Evaluation Metrics. Referring to Golub et al. (1999), we propose precision (PREC) Hossin and Sulaiman (2015), normalized mutual information (NMI) Ferri et al. (2009), and adjusted rand index (ARI) Santos and Embrechts (2009) to evaluate the degree of agreement between the subtyping results obtained by different methods and the true labels.
(2) Cancer Subtype Clustering and Golden-Standard Subtype Classification:
Motivation.
Each specific cancer comprises multiple subtypes. Cancer clustering or classification aims to categorize patients into subgroups based on their multi-omics data. The reason is that while the subtypes may differ in their biochemical levels, they often share the same morphological traits, such as physical structure and form in an organism. However, for most cancer types, subtyping a cancer is still an open question under discussion. Thus, cancer subtyping tasks are typically clustering tasks without ground true labels. Here, the cancer research community has thoroughly analyzed the subtypes of some of the most common cancer types in a previous study. Therefore, we consider these subtypes to contain the true labels and set up a classification task for these subtypes.
[…]

### v2

# MLOmics: Benchmark for Machine Learning on Cancer Multi-Omics Data

@@ Abstract
Framing the investigation of diverse cancers as a machine learning problem has recently shown significant potential in multi-omics analysis and cancer research. Empowering these successful machine learning models are the high-quality training datasets with sufficient data volume and adequate preprocessing. However, while there exist several public data portals including The Cancer Genome Atlas (TCGA) multi-omics initiative or open-bases such as the LinkedOmics, these databases are not off-the-shelf for existing machine learning models. In this paper we propose MLOmics, an open cancer multi-omics benchmark aiming at serving better the development and evaluation of bioinformatics and machine learning models. MLOmics contains 8,314 patient samples covering all 32 cancer types with four omics types, stratified features, and extensive baselines. Complementary support for downstream analysis and bio-knowledge linking are also included to support interdisciplinary analysis.
Background & Summary
Multi-omics analysis has shown great potential to accelerate cancer research. A promising trend consists in framing the investigation of diverse cancers as a machine learning problem, where complex molecular interactions and dysregulations associated with specific tumor cohorts are revealed through integration of multi-omics data into machine learning models. Several impressive achievements have been demonstrated in molecular subtyping [1, 2, 3], disease-gene association prediction [4, 5, 6], and drug discovery [7].
Empowering successful machine learning models are the high-quality training datasets with sufficient data volume and adequate preprocessing. While there exist several public data portals including The Cancer Genome Atlas (TCGA) multi-omics initiative [8] or open-bases such as the LinkedOmics [9], these databases are not off-the-shelf for existing machine learning models. To make these data model-ready, a series of laborious, task-specific processing steps such as metadata review, sample linking, and data cleaning are mandatory. The domain knowledge required, as well as a deep understanding of diverse medical data types and proficiency in bioinformatics tools have become an obstacle for researchers outside of such backgrounds. The gap between the growing body of powerful machine learning models and the absence of well-prepared public data has become a major bottleneck. Currently, some existing research validates proposed machine learning models using inconsistent experimental protocols, including variations in datasets and data processing techniques, and evaluation strategies [10]. These studies could benefit from a fair assessment of extensive baselines on a uniform footing with unified datasets. A benchmark also play a key role in guiding bioinformatics researchers in designing and handpicking the most suitable models.
Figure 1: Schematic workflow of creating the MLOmics. The process starts with collecting patient samples covering 32 cancer types from the TCGA project. All resources in diverse data types and sizes are uniformly integrated and processed to contain data of four omics types. Datasets for benchmark ML tasks were constructed based on the processed data. MLOmics also selected baselines, metrics, and resources to support downstream biological analysis. Overview of the MLOmics. MLOmics provides a user-friendly interface for developing and evaluating machine learning models based on cancer multi-omics data. MLOmics provides datasets in three feature scales for 20 classification, clustering, and omics imputation learning tasks. MLOmics also provides statistical, ML, and DL baselines for each task, which are evaluated by fair metrics. Bio-knowledge database linking with MLOmics. MLOmics provides resources to link with other bio-knowledge databases, enabling integration of outer resources for applications such as ML evaluation, gene-disease association exploration, network inference, and functional analysis.
To meet the growing demand of the community, we introduce MLOmics, an open cancer multi-omics benchmark aiming at serving better the development and evaluation of bioinformatics and machine learning models. MLOmics collected 8,314 patient samples covering all 32 cancer types from the TCGA project. All samples were uniformly processed to contain four omics types: mRNA expression, microRNA expression, DNA methylation, and copy number variations, followed by categorization, protocol verification, feature profiling, transformation, and annotation. Based on the processed data, 20 learning tasks and the associated datasets for pan-cancer analysis, cancer subtypes, and omics imputation were constructed. For each dataset, we provide three feature versions: Original, Aligned, and Top, to support feasible analysis. For example, the Top version contains the most significant features selected via the ANOVA test [11] across all samples to filter out potentially noisy genes. The MLOmics datasets were carefully examined with 6$\sim$10 high citations baseline methods. These baselines were rigorously reproduced and evaluated with various metrics to ensure fair comparison. Complementary resources are included to support basic downstream biological analysis, such as clustering visualization, survival analysis, and Volcano plots. Last but not least, we provide support to interdisciplinary analysis via our locally deployed bio-base resources. Interdisciplinary researchers can retrieve and integrate bio-knowledge of cancer omics studies through resources such as the STRING [12] and KEGG [13]. For instance, exploration of bio-network inference [14] and simulated gene knockouts [15] is supported. In summary, MLOmics provides a user-friendly interface that enables non-experts to easily utilize its open, well-prepared datasets for developing/evaluating machine learning models; conducting interdisciplinary analysis and supporting cancer research and broader biological studies. A detailed overview of the database and its characteristics is provided in Figure. 1.
Methods
Data Collection and Preprocessing
The data were sourced from TCGA via the Genomic Data Commons (GDC) Data Portal [16]. The original resources in TCGA are organized by cancer type, and the omics data for individual patients are scattered across multiple repositories. Therefore, retrieving and collecting omics data require linking samples with metadata and applying different preprocessing protocols. MLOmics employs a unified pipeline that integrates data preprocessing, quality control, and multi-omics assembly for each patient, followed by alignment with their respective cancer types. Specifically, we perform different steps for each omics type as follows:
Transcriptomics (mRNA and miRNA) data: 1. Identifying transcriptomics. We trace the downloaded data using the "experimental_strategy" field in the metadata, marked as "mRNA-Seq" or "miRNA-Seq", then we verify that "data_category" is labeled as "Transcriptome Profiling." 2. Determining experimental platform. We identify the experimental platform from the metadata, such as "platform: Illumina" or "workflow_type: BCGSC miRNA Profiling." 3. Converting gene-level estimates. For data generated by platforms like Illumina Hi-Seq, use the edgeR package [17] to convert scaled gene-level RSEM estimates into FPKM values. 4. Non-Human miRNA filtering. For "miRNA-Seq" data from platforms like Illumina GA and Agilent arrays, we identify and remove non-human miRNA expressions using species annotations from databases such as miRBase [18]. 5. Noise eliminating. We remove features with zero expression in more than 10% of samples or those with undefined value (N/A). 6. Transformation. Finally, we apply logarithmic transformations to obtain the log-converted mRNA and miRNA data.
Genomic (CNV) data: 1. Identifying CNV Alterations. We examine how gene copy-number alterations are recorded in the metadata, using key descriptions such as "Calls made after normal contamination correction and CNV removal using thresholds." 2. Filter Somatic Mutations. We capture only somatic variants by retaining entries marked as "somatic" and filtering out germline mutations. 3. Identify Recurrent Alterations. We use the GAIA package [19] to identify recurrent genomic alterations in the cancer genome, based on raw data representing all aberrant regions from copy-number variation segmentation. 4. Annotate Genomic Regions. We annotate the recurrent aberrant genomic regions using the BiomaRt package [20].
Epigenomic (Methy) data: 1. Identify Methylation Regions. We examine how methylation is defined in metadata to map methylation regions to genes, using key descriptions like "Average methylation (beta-values) of promoters defined as 500bp upstream & 50 downstream of Transcription Start Site (TSS)" or "With coverage >= 20 in 70% of the tumor samples" 2. Normalize Methylation Data. We perform median-centering normalization to adjust for systematic biases and technical variations across samples, using the R package limma [21]. 3. Select Promoters with Minimum Methylation. For genes with multiple promoters, we select the promoter with the lowest methylation levels in normal tissues.
After processing the omics sources, the data are annotated with unified gene IDs to resolve variations in naming conventions caused by the difference in sequencing methods or reference standards [22]. Then, the omics data are aligned across multiple sources based on their corresponding sample IDs. Finally, the data files are organized by cancer type for further dataset construction.
Datasets Construction
MLOmics reorganizes the collected and processed data resources into different feature versions tailored to various machine learning tasks. For each task, MLOmics provides several baselines, evaluation metrics, and the ability to link with biological databases such as STRING and KEGG for further biological analysis of different machine learning models.
Feature Processing.
Machine learning models require tabular data with a the same number of features across samples. In addition to the Original feature scale that contains a full set of genes (variations included) directly extracted from the collected omics files, MLOmics provides two other well-processed feature scales: Aligned and Top. The former scale filters out non-overlapping genes and selects the genes shared across different cancer types; and the latter identifies the most significant features. Specifically, the following steps are performed for each scale:
Aligned: 1. we resolve the mismatches in gene naming formats such as ensuring compatibility between cancers that use different reference genomes. 2. we identify the intersection of feature lists across datasets to ensure all selected features are present in different cancers. 3. we conduct z-score feature normalization.
Top: 1. we perform multi-class ANOVA [11] to identify genes with significant variance across multiple cancer types. 2. we perform multiple testing using the Benjamini-Hochberg (BH) [23] correction to control the false discovery rate (FDR) [24]. 3. the features are ranked by the adjusted $p$-values $p<0.05$ or by the user-specified scales). 4. we conduct z-score feature normalization which reduces the presence of non-significant genes across cancers and this could be beneficial for biomarker studies.
20 Task-ready Datasets with Baselines and Metrics.
We provide 20 off-the-shelf datasets ready for machine learning models ranging from pan-cancer/cancer subtype classification, subtype clustering to omics data imputation. We also include well recognized baselines that leveraged classical statistical approaches and machine/deep learning methods as well as metrics for standard evaluation.
Pan-cancer and golden-standard cancer subtype classification. Pan-cancer classification aims to identify each patient’s specific cancer type. Moreover, a cancer typically comprises multiple subtypes that differ in their biochemical profiles. Some subtypes have been well-studied in prior research and widely accepted as the golden standard. We re-label patient samples to support subtyping evaluation. These two classification tasks potentially improve cancer early diagnostic accuracy and treatment outcomes.
Datasets: MLOmics provides six labeled datasets: one pan-cancer dataset and five gold-standard subtype datasets (GS-COAD, GS-BRCA, GS-GBM, GS-LGG, and GS-OV).
Baselines: we opt for the following classical classification methods as baselines: XGBoost [25], Support Vector Machines (SVM) [26], Random Forest (RF) [27], and Logistic Regression (LR) [28]. Additionally, we include six popular, open-sourced deep learning methods: Subtype-GAN [29], DCAP [30], XOmiVAE [31], CustOmics [32], and DeepCC [33].
Metrics: For classification evaluation, we opt for precision (Pre), recall (Re), and F1-score (F1). Since clustering is the primary focus of the subtyping task, due to the limited sample size (typically <100), we provide normalized mutual information (NMI) and adjusted rand index (ARI) to evaluate the agreement between the clustering results obtained by different methods and the true labels.
Cancer Subtype Clustering. Cancer subtyping remains an open question under fierce debate for most cancers, especially rare types. Numerous studies propose various clustering methods to identify distinct groups by identifying different clusters to support downstream evaluation and discovery of new subtypes.
Datasets: MLOmics provides nine unlabeled rare cancer datasets (ACC, KIRP, KIRC, LIHC, LUAD, LUSC, PRAD, THCA, and THYM) for this learning task.
Baselines: In addition to the aforementioned Subtype-GAN, DCAP, MAUI, XOmiVAE, we also include six clustering methods: Similarity Network Fusion (SNF) [34], Neighborhood-based Multi-Omics clustering (NEMO) [35], Cancer Integration via Multi-kernel Learning (CIMLR) [36], iClusterBayes [37], moCluster [38], and MCluster-VAEs [39].
Metrics: To evaluate the goodness of clustering we opt for the classic Silhouette coefficient (SIL) [40] and log-rank test $p$-value on survival time (LPS) [41].
Omics Data Imputation. In addition to classification and clustering, we also provide a data imputation task focusing on imputing multi-omics cancer data. The collected omics data typically have missing values due to experimental limitations, technical errors, or inherent variability. The imputation process is crucial for ensuring the integrity and usability of TCGA omics data [42].
Datasets. MLOmics provides three omics datasets with missing values (GS-BRCA, GS-COAD, GS-GBM). Given a full dataset as a matrix $D\in\mathbb{R}^{n\times m}$, we follow prior works [43, 42] to generate a mask matrix $M\in\{0,1\}^{n\times m}$ uniformly at random with the probability of missing $P(M_{ij}=0)=r_{\text{miss}}$, and the probability of retaining $P(M_{ij}=1)=1-r_{\text{miss}}$. The final data matrix is obtained by multiplying element-wise the data matrix $D$ with the mask $M$. The missing level $r_{\text{miss}}$ is selected from $[0.3,0.5,0.7]$.
Baselines. We opt for seven well-recognized methods for imputing missing values, including: Mean Imputation that imputes the missing entry with mean values of entries around it (Mean); K-Nearest Neighbors (KNN) that imputes the missing value with the weighted Euclidean K nearest neighbors; Multivariate Imputation by Chained Equations (MICE) that performs multiple regression to model each missing value conditioned on non-missing values [44]; Iterative SVD (iSVD) that imputes by iterative low-rank SVD decomposition [45]; Spectral Regularization Algorithm (Spectral) that also employs SVD but with a soft threshold and the nuclear norm regularization [46]; Generative Adversarial Imputation Nets (GAIN) that proposes to distinguish between fake and true missing patterns by the generator-discriminator architecture [43]; Graph Neural Network for Tabular Data (GRAPE) that utilizes the graph networks to impute based on learned information from columns and rows of the data matrix [42].
Metrics. We use the Mean Squared Error (MSE) between the unmasked entries ($M_{ij}=1$) as the training loss to let the model predict the actual value. During test, the masked missing values are used for evaluating the model performance ($M_{ij}=0$).
MLOmics will be continuously updated with baselines and evaluations on the defined learning tasks. Detailed description of the baselines and metrics is provided in the Supplementary Material.
Biological Database Linking and Downstream Analysis
A rising trend in multi-omics analysis is to integrate multi-omics data (non-network data) with biological networks to better understand complex functions on the gene or protein level. MLOmics provides offline linking resources for well-established databases such as STRING [12] and KEGG [13]. STRING offers network structural information about genes, while KEGG provides systematic functional information on various biological networks.
To integrate omics data into biological networks, omics features (e.g., genes) are first mapped to STRING and KEGG network nodes (e.g., gene products) using standardized identifiers, e.g. unified gene IDs. MLOmics provides mapping files to extract shared gene IDs and standardize inconsistent file formats, creating a unified feature list. This feature list includes shared gene IDs along with corresponding omics features or measurements. Once the omics features are successfully mapped, they can be incorporated into networks by assigning weights to nodes or edges based on the data e.g. expression levels. Additionally, MLOmics provides a suite of biological analysis tools to evaluate the significance of results generated by different machine learning models. These tools include survival analysis [47], gene differential expression (GDE) analysis [48], featuring volcano plots and simulated knockout analyses at the single-gene level, and KEGG pathway analysis [49], which operates at the gene function set level. These resources are designed to support researchers in conducting classic and reliable biological validations of their data analysis frameworks.
Data Records
As shown in Fig. 2, MLOmics framework consists of three major components: (1) main datasets that includes all cancer multi-omics datasets for various tasks; (2) baselines and metrics that provides the source code for baseline models and performance metrics and (3) downstream analysis tools and reources linking that provides tools for further analysis and links to additional biological resources.
Figure 2: Schematic of MLOmics resources structure. MLOmics framework consists of three major components. The Main Datasets file includes all cancer multi-omics datasets for various tasks. The Baseline and Metrics file provides the source code for baseline models and performance metrics. The Downstream Analysis Tools and Resources Linking file offers sources for further analysis and links to additional biological resources.
Figure 3: The experiment results and downstream analysis of applying machine learning baselines based on MLOmics datasets. (a) The PREC bar plot for each baseline method across all datasets. (b) The SIL heatmap for each baseline method across all datasets. (c) The box plot for each baseline method on three imputation datasets. (d-e) The downstream analysis results based on the clustering outcomes of XOmiVAE applied to the BRCA patient clustering datasets.
Main Datasets
[…]